Wallis Company manufactures only one product and uses a standard cost system. The company uses a predetermined plantwide overhead rate that relies on direct labor-hours as the allocation base. All of the company's manufacturing overhead costs are fixed—it does not incur any variable manufacturing overhead costs. The predetermined overhead rate is based on a cost formula that estimated $2,919,000 of fixed manufacturing overhead for an estimated allocation base of 291,900 direct labor-hours. Wallis does not maintain any beginning or ending work in process inventory.

The company’s beginning balance sheet is as follows:

Wallis Company
Balance Sheet
1/1/XX
(dollars in thousands)
Assets
Cash $ 890
Raw materials inventory 340
Finished goods inventory 460
Property, plant, and equipment, net 10,400
Total assets $ 12,090
Liabilities and Equity
Retained earnings $ 12,090
Total liabilities and equity $ 12,090

The company’s standard cost card for its only product is as follows:

Inputs (1)
Standard
Quantity
or Hours
(2)
Standard
Price
or Rate
Standard
Cost
(1) × (2)
Direct materials 2 pounds $ 33.80 per pound $ 67.60
Direct labor 3.00 hours $ 14.00 per hour 42.00
Fixed manufacturing overhead 3.00 hours $ 10.00 per hour 30.00
Total standard cost per unit $ 139.60

During the year Wallis completed the following transactions:

  1. Purchased (with cash) 239,500 pounds of raw material at a price of $31.40 per pound.
  2. Added 219,750 pounds of raw material to work in process to produce 96,900 units.
  3. Assigned direct labor costs to work in process. The direct laborers (who were paid in cash) worked 248,800 hours at an average cost of $16.00 per hour to manufacture 96,900 units.
  4. Applied fixed overhead to work in process inventory using the predetermined overhead rate multiplied by the number of direct labor-hours allowed to manufacture 96,900 units. Actual fixed overhead costs for the year were $2,749,500. Of this total, $1,359,000 related to items such as insurance, utilities, and salaried indirect laborers that were all paid in cash and $1,390,500 related to depreciation of equipment.
  5. Transferred 96,900 units from work in process to finished goods.
  6. Sold (for cash) 93,900 units to customers at a price of $170 per unit.
  7. Transferred the standard cost associated with the 93,900 units sold from finished goods to cost of goods sold.
  8. Paid $2,129,500 of selling and administrative expenses.
  9. Closed all standard cost variances to cost of goods sold.

Required:

1. Compute all direct materials, direct labor, and fixed overhead variances for the year.

2. Record transactions a through i for Wallis Company.

3. Compute the ending balances for Wallis Company’s balance sheet.

4. Prepare Wallis Company’s income statement for the year.

INTERESTED IN QUESTION 2,3 and 4
Answer #1

Part 1

Materials price variance = AQ*(AC-SC) = 239500*(31.40-33.80) = $574800 F

Materials quantity variance = SC*(AQ-SQ) = 33.80*(219750-(96900*2)) = $877110 U

Labor rate variance = AH*(AR-SR) = 248800*(16-14) = $497600 U

Labor efficiency variance = SR*(AH-SH) = 14*(248800-(96900*3)) = $586600 F

Budget variance = actual FOH – Budgeted FOH = 2749500-2919000 = 169500 F

Volume variance = Budgeted FOH – Applied FOH = 2919000-(96900*3*10) = 12000 U

Part 2 & 3

Wallis company

Transaction analysis

For the year ended 12/31/XX

(dollars in thousands)

Cash

Raw materials

Work-in-progress

Finished goods

PP&E (net)

=

Material price variance

Material quantity variance

Labor rate variance

Labor efficiency variance

Fixed overhead budget variance

Fixed overhead volume variance

Retainer earnings

1/1

890

340

460

10400

12090

a.

(7520)

8095

575

b.

(7428)

6550

(878)

c.

(3981)

4070

(498)

587

d.

(1359)

2907

(1390)

170

(12)

e.

(13527)

13527

f.

15963

15963

g.

(13108)

(13108)

h.

(2130)

(2130)

i.

(575)

878

498

(587)

(170)

12

(56)

1863

1007

0

879

9010

0

0

0

0

0

0

12759

96900*139.60 = 13527240

93900*170 =15963000

93900*139.60 = 13108440

Part 4

Income statement

Sales

15963

Cost of goods sold (standard)

13108

Total variance adjustments

56

Cost of goods sold (actual)

13164

Gross profit (actual)

2799

Selling and administrative expenses

2130

Net income

$699